These were accepted with minor corrections. Typo to be fixed.
The best solution for accessing e.g. printer web interface was considered to be tunnelling/authenticated proxy. Continue to email Ian with suggested tools.
We haven't seen as many problems recently. Alastair will check with Angus about how we should communicate with COs within IS and other Schools. Alison will also email ITPF asking if a) it's a problem for anyone else and b) what methods other COs use. Alison will also put together a list of all the 'hacky' fixes we have in place.
Alison will email COs about nag scripts to make a decision on which we feel are required.
Subsequent to the meeting, Ian noted that venom was still in the AT rack. It has now been removed.
Done - but no tickets have appeared in the queue so Alison will create a test ticket.
We do have on-site cover. Services to look through paperwork for specific case in question.
Alastair has spoken to Mike. Nothing happened on Thu 4th but they will be turned on again on Friday 19th.
Toby has been in touch with Jeff Altman and submitted a bug report to secure-endpoints, who distribute heimdal and NIM for windows. In summary, we should recommend that our users install MIT for Kerberos but note that it has stability issues.
Alastair has also spoken to procurement suggesting default should be 64-bit.
To be removed from here as being tracked at CEG. Will use git and MPU will create a project for it.
Completed - but Alison will check with Stephen.
Completed - on all devel machines - yum2pkglist
Ian suggested that this should be dropped from the list.
Blog articles completed:
Report from Computing Executive Group
Alastair noted that the first of the T2 project prioritisation meetings was held on Monday 8th.
Reports from units
The only time that was considered unsuitable for the reboots was between 7pm and 10pm.
George reminded everyone to set their mh0 interfaces to be the bonding primary and we should use LCFG_OPTIONS_ETHERBOND_PRIMARY rather than setting resources explicitly.
It was suggested that the fix mentioned should be implemented on physical machines only and not virtual machines.
A general point cropped up while discussing the fglrx/Catalyst/ATI/AMD Driver, namely we should update our documentation to include a paragraph explaining that DICE machines will require a re-install if left off for long period of time e.g. more than 2 weeks. A ticket has subsequently been added to the documentation queue - RT61644.
Iain and Graham talked through the fix that they have. Iain hopes to have the sl6 server ready later this week/early next week.
Group 1: Craig, Ian, Iain, Jennifer, Richard has worked one day so far.
Group 2: Neil, Alastair, Toby, Ross has finished.
Group 3: Roger, Chris, George, Graham, Lindsey has finished.
Group 4: Gordon, Stephen, Tim, Alison, Carol has worked one day so far.
It was confirmed that the first free Wednesday in April (the 17th) would be used for groups to give their final reports.
Topics for discussion
AOCBNeil suggested drinks on Friday 26th. Venue to be arranged later.
The next meeting will be on Wednesday 24th April at 10am in the Turing Room, chaired by Alison.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh